### v1

# Targeted Error Correction in Knowledge Distillation: Small Language Models Surpass GPT

Abstract
We introduce an Analyze–Revise–Finetune (ARF) pipeline that enables smaller open-source language models (LLMs) to surpass substantially larger proprietary models in customer service summarization tasks. The pipeline first analyzes and categorizes common errors in summaries produced by a teacher model (GPT-3.5), then performs a targeted revision using a compact editor model (Llama 3.1 70B) to generate high-quality, refined training data. Fine-tuning a smaller student model (Llama 3.1 8B) on this refined data resulted in superior summarization performance compared to GPT-3.5. The ARF pipeline improves cost efficiency and data privacy while maintaining competitive accuracy, illustrating a generalizable framework for enhancing open-source LLMs across diverse downstream applications.
Keywords Automatic error correction $\cdot$ LLM fine-tuning $\cdot$ Summarization
1 Introduction
We are in a new era of large language models (LLMs), with numerous products widely deploying these powerful models. This widespread deployment necessitates continuous improvement and monitoring of model quality, particularly concerning factors such as output quality, privacy, and cost. Proprietary LLMs, like GPT, are commonly employed in production environments due to their high-quality outputs achievable through prompt engineering. However, reliance on such large proprietary models carries significant drawbacks, including high costs and the potential risk of price increases. Additionally, off-site hosting of these models introduces privacy and data security concerns, as sensitive data must be transmitted to external servers. Consequently, utilizing smaller, open-source LLMs tailored through task-specific fine-tuning is becoming increasingly preferable, offering reduced inference costs and mitigating privacy risks through on-premises deployment.
Task-specific fine-tuning substantially improves the quality of text generation in smaller LLMs. However, creating a sufficiently large dataset for training can become a critical bottleneck, especially when such data must be generated manually. To address this issue, recent methodologies have utilized pseudo-labeling strategies, where text generated by larger (teacher) LLMs serves as training labels for smaller (student) LLMs[1, 2, 3]. These approaches enable student models to achieve performance close to that of teacher models. However, they also inadvertently propagate any errors present in teacher-generated texts, as all outputs from the teacher model are accepted as correct.
To enhance the quality of pseudo-labels produced by teacher models, recent research has explored self-improvement methods for text generation [4, 5, 6]. These methods typically involve iterative critic-and-revision cycles, in which an LLM evaluates its own output and subsequently revises the text based on its self-critique. Although this iterative approach results in improved final outputs, it incurs substantial computational and inference costs, particularly when applied across extensive training datasets. Moreover, the critic and generation tasks typically rely on the same advanced LLM, increasing the overall inference expense.
In this paper, we introduce an "Analyze-Revise-Finetune" (ARF) pipeline (illustrated in Figure 1) designed to enhance the performance of smaller, open-source LLMs (student models) so they surpass the capabilities of larger, proprietary teacher models. ARF comprises three primary steps: 1) systematic identification of prevalent error patterns in auto-generated texts, 2) targeted error correction using a separate editor LLM, and 3) leveraging the corrected texts to fine-tune a smaller open-source student LLM. Unlike previous studies that either exclusively depend on teacher-generated pseudo-labels or iterative self-refinement, ARF incorporates a targeted error correction step based on an error-focused evaluation scheme that facilitates a systematic analysis of teacher-generated texts to pinpoint prevalent errors in the training data. By specifically targeting these common errors for correction, we remove the need for repetitive critic-and-revision cycles. Furthermore, targeted error correction is inherently simpler than full critic-revision, allowing the editor LLM to be smaller and more cost-effective than the teacher model. By fine-tuning the student LLM using the revised training data, the student model learns effectively from the teacher without inheriting its frequent errors, unlike methods that uncritically adopt all teacher-generated pseudo-labels.
We demonstrate the effectiveness of ARF pipeline in a customer service interaction summarization task currently served by GPT-3.5. Initially, we define comprehensive evaluation criteria to systematically identify and classify specific error patterns present in GPT-3.5-generated summaries. Utilizing these criteria, we identify common errors prevalent across summaries. Subsequently, we employ a smaller editor model (Llama 3.1 70B) to specifically revise these identified errors, thus creating a refined training dataset. Finally, we fine-tune a smaller open-source model (Llama 3.1 8B) on this revised dataset, demonstrating that our fine-tuned model successfully replicates GPT-3.5’s functionality while significantly reducing frequent errors, ultimately outperforming GPT-3.5 in summary generation.
Figure 1: "Analyze-Revise-Finetune" pipeline
2 Related Work
Many studies leverage larger models ("teacher" models) to generate pseudo-labels for training smaller models ("student" models). Hinton et al. [1] introduced a knowledge distillation method where the teacher model’s output probabilities serve as soft targets for the student model training. This approach enables the student to capture nuanced predictions from the teacher, effectively consolidating knowledge from an ensemble of models into a single model. Gu et al. [2] extended knowledge distillation to large language models (LLMs), employing reverse Kullback-Leibler divergence (KLD) instead of the conventional forward KLD, comparing the student LLM’s output distribution directly with that of the teacher. Sanh et al. [3] developed DistilBERT, which achieves 97% of BERT’s language understanding capability while being 40% smaller by minimizing the KL divergence between student and teacher softmax outputs. Unlike these approaches, which typically treat all teacher-generated outputs equally without any error analysis or correction, ARF provides a focused approach where common errors in teacher-generated texts are explicitly identified and corrected, thereby ensuring higher-quality pseudo-labels.
Recent research has investigated iterative refinement techniques involving LLMs, where a model evaluates and improves its own generated outputs. "Self-Refine"[4] is a framework that allows an LLM to iteratively critique and refine its responses autonomously without external supervision or additional training data. Similarly, LLM-AUGMENTER[5] addresses hallucinations in LLM outputs by integrating external knowledge and iterative prompt revision based on feedback from utility functions. Bai et al. [6] proposed Constitutional AI, a self-supervised framework that utilizes guiding principles (a "constitution") to critique and iteratively revise an LLM’s outputs, ultimately using these revised outputs to train a safer model. In contrast to these iterative self-refinement methods, ARF emphasizes targeted correction of the most common errors identified through systematic analysis. Additionally, while previous methods generally rely on a single LLM for both output generation and self-correction, ARF separates initial generation (performed by teacher LLM) from subsequent error correction (performed by editor LLM). This separation enables the use of a smaller and more cost-efficient editor model, achieving high-quality revisions with lower inference costs.
3 Background: Customer Service Interactions Summarization with GPT3.5
In 2024, eBay deployed an LLM-based summarization system aimed at enhancing the productivity of customer service agents (referred to as teammates). To effectively assist customers, teammates must first understand the customer’s issue or request by reviewing any previous interactions the customer had with eBay. The primary role of the summarization system is to succinctly summarize these prior interactions, highlighting essential information while removing any superfluous details, enabling teammates to efficiently and promptly take appropriate next steps.
This summarization system processes textual input from customer interactions originating from four distinct communication channels: bot-customer chat, teammate-customer chat, emails, and webforms. Before summarization, the system automatically redacts personal information from the input content. The sanitized text is then sent to GPT-3.5, which generates concise summaries presented as bullet points within the Service Desktop interface. To facilitate continuous improvement, a feedback mechanism allowing teammates to give a thumbs-up or thumbs-down to indicate summary quality is integrated into the Service Desktop. Since its initial launch in February 2024, the summarization system has maintained a thumbs-down feedback rate below 5%.
We fine-tune a compact open-source LLM for customer service summarization with the goal of replacing GPT-3.5 in production. This approach substantially reduces inference costs and mitigates privacy risks inherent to proprietary models. Our experiments focus on two communication channels—bot–customer chats (BotChat) and webforms (WebForm)—which account for most high-volume, text-based interactions in the production environment.
4 Error-Focused Summary Evaluation Scheme
To identify opportunities for further enhancing the quality of the production summarizer and to support AI governance efforts, we retrieved a representative sample of input content and corresponding summaries from production data. These samples were then manually reviewed by a team of subject matter experts and the AI governance team. We collaborated closely with this team to develop a structured summary evaluation scheme aimed at ensuring consistency in the review process and facilitating systematic identification of recurring issues and potential areas for improvement. This evaluation scheme is centered around the identification of specific "errors" within summaries, categorized based on predefined error types derived from an initial data analysis.
Errors are grouped into seven primary categories: Content, Entities, Data Elements, Customer Type, Unnecessary Information, Inferred Sentiment, and Language. Each category is further divided into detailed sub-labels specifying the precise nature of errors. The Content category covers fundamental summary issues such as missing critical details, inaccuracies, or fabricated content. The Entities category addresses problems related to missing or inaccurately represented key identifiers, including item numbers, names, order numbers, and other transaction-related details. Data Elements encompasses inaccuracies in structured information, such as dates, that are not captured by the Entities category. Customer Type focuses on incorrectly identifying or omitting the customer’s role (buyer or seller), a specific error relevant only to WebForm data. Unnecessary Information flags the inclusion of irrelevant or redundant details, including unnecessary bot interactions, courteous engagements, or repeated information. Inferred Sentiment targets unwarranted emotional inferences about the customer that are unsupported by the original interaction. Lastly, Language captures errors associated with incorrect or missing language identification, as well as translation inaccuracies. This comprehensive categorization framework supports a consistent, precise, and detailed evaluation of summary quality. A complete taxonomy of summary error types is available in Table 5 in the appendix.
Summaries are rated on a 1-to-5 scale according to the number and severity of errors identified. A rating of 1 indicates numerous or critical errors significantly impacting the overall understanding of the situation and the subsequent customer service. A rating of 2 denotes summaries containing one or two errors. A summary receives a rating of 3 if it contains only a minor error that does not affect the overall understanding. Summaries rated 4 are considered acceptable, with no distinct errors but potentially lacking perfect conciseness. Finally, a rating of 5 represents summaries that are flawless, containing no errors or issues.
We utilize this evaluation scheme to assess the quality of summaries generated by various LLMs and to compare the performance of different fine-tuned models (Section 6). Additionally, we use the predefined error types in the scheme to systematically analyze GPT-3.5-generated summaries and identify the most common errors (Section 5.2).
5 LLM Fine-Tuning via Targeted Error-Correction
5.1 Data
Due to the presence of personally identifiable information (PII) in customer interaction logs, we developed a secure data retrieval pipeline that downloads production data samples from the RNPCI zone and anonymizes all PII by substituting them with artificial fake values. Using this pipeline, we collected 10,000 samples per communication channel to serve as training data. Specifically for WebForm data, additional preprocessing steps were applied to parse the XML structure and extract only the information pertinent to customer service interactions. Furthermore, we retrieved separate test and development sets comprising 200 samples each per channel. Lastly, smaller datasets consisting of 100 BotChat and 68 WebForm samples were obtained to facilitate detailed analysis of prevalent error types in production-generated summaries.
5.2 Systematic error analysis
To identify the most frequent error types occurring in summaries produced by GPT-3.5, a team of subject matter experts reviewed production summaries—specifically 100 BotChat samples and 68 WebForm samples. Each summary was evaluated for errors based on the predefined error types outlined in our summary evaluation scheme (Table 5). After completing the evaluation, we determined the most common error types present in the summaries from each channel. Table 1 summarizes these key error types.
Table 1: Major error types across different channels
Bot Chat WebForm
Error Type Count Error Type Count
unn_content_requests_agent 30 (38.0%) unn_content_redundant 14 (29.2%)
sentiment_inferred_frustrated 19 (24.0%) customer_type_inaccurate 7 (14.6%)
content_missing 13 (16.5%) unn_content_webform_email_copy 4 (8.3%)
nothing_to_summarize 6 (7.6%) entity_missing_item_number 3 (6.3%)
sentiment_inferred_not_frustrated 5 (6.3%) content_inaccurate 3 (6.3%)
all errors 79 (100%) all errors 48 (100%)
5.3 Automatic error-correction
From the error types identified through systematic analysis (Section 5.2), we selected four frequent and manageable errors for automated correction. For BotChat, we targeted the errors unn_content_requests_agentunn\_content\_requests\_agent, referring to unnecessary mentions in summaries of customers requesting human assistance, and sentiment_inferred_frustratedsentiment\_inferred\_frustrated, referring to inferred customer frustration not explicitly stated by the customer. For WebForm summaries, we selected the errors unn_content_redundantunn\_content\_redundant, denoting unnecessary redundant information, and unn_content_webform_email_copyunn\_content\_webform\_email\_copy, indicating the unnecessary inclusion in summaries of customer requests for email copies of their submitted webforms.
We employed Llama 3.1 70B as the editor LLM for performing these error corrections. Being an open-source model, Llama 3.1 70B provides reduced privacy risks. Correction prompts designed for each error type are depicted in Figure 2. Specifically, the prompt addressing sentiment_inferred_frustratedsentiment\_inferred\_frustrated was designed broadly to remove all references to sentiment rather than just inferred frustration. This approach does not negatively impact summary quality since sentiment information is considered non-essential according to our evaluation criteria (Section 4).
We processed all 20,000 training samples across both channels using a cascading sequence of error-correction prompts. For BotChat summaries, corrections were first applied to remove sentiment_inferred_frustratedsentiment\_inferred\_frustrated, followed by unn_content_requests_agentunn\_content\_requests\_agent. For WebForm summaries, corrections began with unn_content_webform_email_copyunn\_content\_webform\_email\_copy. Subsequently, the prompt addressing unn_content_redundantunn\_content\_redundant was applied to both channels, under the assumption that increased conciseness would benefit both datasets. We denote the dataset prior to applying the unn_content_redundantunn\_content\_redundant fix as revision version 1 (r1), and the dataset after applying this correction as revision version 2 (r2).
Prompt to remove unn_content_requests_agentunn\_content\_requests\_agent error from BotChat summaries Read the following summary of a chat transcript between an eBay customer and an automatic service agent. Rewrite the summary to remove any information about the customer wanting or requesting to connect with a human agent. If there’s no other content in the summary other than customer wanting or requesting to connect with a human agent, say "<ul><li>nothing to summarize<li></ul>". Keep the original format of the summary that is unordered list items in HTML. Do not add anything else other that the revised summary.
Prompt to remove sentiment_inferred_frustratedsentiment\_inferred\_frustrated error from BotChat summaries Read the following summary of a chat transcript between an eBay customer and an automatic service agent. Rewrite the summary to remove any information about the customer’s emotion such as being frustrated, any information about the reason of the emotion, and any information about how the customer expressed their emotion. If there’s no other content in the summary other than customer’s emition, say "<ul><li>nothing to summarize<li></ul>". Keep the original format of the summary that is unordered list items in HTML. Do not add anything else other that the revised summary.
Prompt to remove unn_content_webform_email_copyunn\_content\_webform\_email\_copy error from WebForm summaries Read the following summary of a WebForm submitted by an eBay customer. Rewrite the summary to remove any information about the customer wanting or requesting a copy of the communication to their email. Do not add anything else other that the revised summary.
Prompt to remove unn_content_redundantunn\_content\_redundant for both BotChat and WebForm summaries Read the following summary. Rewrite the summary to remove any information that is repeatedly stated. Do not remove any important information such as Item Ids, Return Ids, Case Ids, etc or customer’s role as a buyer or a seller. Keep the original format of the summary that is unordered list items in HTML. Do not add anything else other that the revised summary.
Figure 2: Prompts for summary error correction
5.4 LLM fine-tuning
We evaluated multiple Llama and Lilium models of different sizes as student LLMs. Each model was fine-tuned independently using three dataset versions: the original dataset (org), revision version 1 (r1), and revision version 2 (r2). Instruction-tuning samples were generated using the same summary generation prompts currently utilized by the production GPT-3.5 model. Fine-tuning was performed with Low-Rank Adaptation (LoRA) [7], applying parameters $alpha=16$, $dropout=0.1$, and $r=8$ across all linear modules.
Fine-tuning of the Lilium 2 mix 7B and Lilium 2 13B models was conducted using Athena, while other models utilized between 4 and 8 A100 GPUs depending on model size. During inference, we used the same prompts as those employed in the production environment for all experiments.
6 Evaluation
Summaries produced by the fine-tuned student LLMs were evaluated using our predefined summary evaluation scheme (Section 4) and assigned ratings on a 1-to-5 scale. Since manual evaluation of summaries can be time-intensive and could potentially slow down experimental workflows, we developed an automatic summary evaluation system leveraging the LLMs-as-a-Judge framework [8, 9].
Specifically, we used GPT-4 Turbo as the judge LLM and performed prompt tuning separately for each communication channel to output ratings on a 1-to-5 scale, based on the original content and corresponding generated summaries. Our automatic evaluation system demonstrated strong correlation with human ratings when benchmarked against other advanced LLM-as-judge methodologies [8, 9]. For the BotChat channel, the correlation was measured with a Spearman’s $\rho$[10] of 0.6663 and Kendall’s $\tau$[11] of 0.6005. For the WebForm channel, Spearman’s $\rho$ was 0.5674, and Kendall’s $\tau$ was 0.6364.
Our primary evaluation metric is the mean score of automatic ratings provided by our evaluators on the test dataset.
7 Results
We first evaluated the effectiveness of the "Analyze-Revise-Finetune" pipeline across various student LLMs. Table 2 presents the summarization performance for each model across different training conditions. Each model was assessed under four distinct scenarios: 1) without fine-tuning (out-of-box), 2) fine-tuned on the original training data (org), 3) fine-tuned on revision version 1 of the training data (r1), and 4) fine-tuned on revision version 2 of the training data (r2). To clearly illustrate the impact of revised training data, we include the score differences compared to the original dataset (org) for both r1 and r2 ($\Delta$ over org.).
Table 2: Summarization performance across different training data sets and student LLMs. Mean of auto-evaluator rating (1-5 scale) is given. Superscript * marks performance better than teacher LLM (GPT-3.5). Best scores for each channel are bold-faced.
Bot Chat WebForm
Mean auto-rating $\Delta$ over org. Mean auto-rating $\Delta$ over org.
GPT3.5 (out-of-box) 4.05 4.29
Llama 3.1 8B (out-of-box) 2.34 3.32
Llama 3.1 8B (org) 4.14* 4.32*
Llama 3.1 8B (r1) 4.325* 0.185 4.405* 0.085
Llama 3.1 8B (r2) 4.315* 0.175 4.125 -0.195
Llama 3.1 70B (out-of-box) 4.28* 4.39*
Llama 3.1 70B (org) 3.96 4.365*
Llama 3.1 70B (r1) 4.315* 0.355 4.42* 0.055
Llama 3.1 70B (r2) 4.3* 0.34 4.22 -0.145
Lilium 2 7B (out-of-box) 1.145 1.975
Lilium 2 7B (org) 3.975 4.125
Lilium 2 7B (r1) 4.28* 0.305 4.295* 0.17
Lilium 2 7B (r2) 4.2* 0.225 4.185 0.06
Lilium 2 mix 7B (out-of-box) 1.015 1.47
Lilium 2 mix 7B (org) 3.825 4.125
Lilium 2 mix 7B (r1) 4.205* 0.38 4.11 -0.015
Lilium 2 mix 7B (r2) 4.16* 0.335 3.92 -0.205
Lilium 2 13B (out-of-box) 1.21 2.22
Lilium 2 13B (org) 3.755 3.995
Lilium 2 13B (r1) 4.075* 0.32 4.065 0.07
Lilium 2 13B (r2) 4.045* 0.29 3.8 -0.195
Lilium 2 mix 45B (out-of-box) 1.01 1.04
Lilium 2 mix 45B (org) 3.89 1.38
Lilium 2 mix 45B (r1) 4.225* 0.335 1.185 -0.195
Lilium 2 mix 45B (r2) 3.95 0.06 1.145 -0.235
Overall, the top-performing models were Llama 3.1 8B for the BotChat channel and Llama 3.1 70B for the WebForm channel, each fine-tuned using the r1 training set. Notably, these student models outperformed the GPT-3.5. This clearly demonstrates the effectiveness of the "Analyze-Revise-Finetune" method in enhancing smaller models to surpass larger, proprietary teacher models through targeted data revisions.
For the BotChat summarization task, all evaluated models exhibited performance improvements using both r1 and r2 revised training datasets, underscoring the benefit of our pipeline. Importantly, nearly all models fine-tuned on the revised datasets—except Lilium 2 mix 45B (r2)—exceeded the performance of GPT-3.5.This demonstrates that our "Analyze-Revise-Finetune" approach can effectively overcome the typical performance limitations associated with traditional knowledge distillation methods. In contrast, models fine-tuned solely on the original training set did not surpass GPT-3.5, with the single exception of Llama 3.1 8B.
In the WebForm summarization task, Llama 3.1 8B, Llama 3.1 70B, Lilium 2 7B, and Lilium 2 13B benefitted notably from the r1 training dataset, again surpassing GPT-3.5 and highlighting the effectiveness of fine-tuning through targeted error-correction for specific tasks. Interestingly, the two Lilium mixture-of-expert models (Lilium 2 mix 7B and Lilium 2 mix 45B) exhibited negative performance impacts when using the r1 dataset; further investigation into this phenomenon is left for future research.
However, for the WebForm summarization task, the r2 training dataset resulted in performance degradation for most student LLMs, except Lilium 2 7B. The reduced performance can be attributed to the quality of the revised summaries. As illustrated in Table 3, the revision success rate varied by error type. The revision addressing unn_content_webform_email_copyunn\_content\_webform\_email\_copy had a high success rate of 97%, generating the r1 dataset. In contrast, the revision targeting unn_content_redundantunn\_content\_redundant, creating the r2 dataset, had a success rate of only 61%. This lower revision quality also explains the smaller performance improvements seen in BotChat summaries when using the r2 dataset compared to the performance improvement using r1 dataset.
Table 3: Summary revision success rate per error type
Error type Bot Chat WebForm Overall
unn_content_requests_agent 94% - 94%
sentiment_inferred_frustrated 92% - 92%
unn_content_webform_email_copy - 97% 97%
unn_content_redundant 78% 61% 70%
Additionally, we explored how the sequence or order in which training data from different channels is provided affects summarization performance. Table 4 presents the summarization results for Llama 3.1 8B trained on the r1 dataset with various training sequences. Initially, we created two models, each trained exclusively on one channel: BotChat-only and WebForm-only. Then, these models were sequentially fine-tuned on data from the other channel, resulting in two additional models: one trained first on BotChat and then on WebForm (BotChat→WebForm), and another trained first on WebForm and then on BotChat (WebForm→BotChat). Finally, we trained a model using data from both channels simultaneously (BotChat+WebForm), which represents the default configuration for all other experiments discussed in this paper.
Table 4: Summarization performance of Llama 3.1 8B trained on r1 with different training sequences. Mean of auto-evaluator rating (1-5 scale) is given. Best scores for each channel are bold-faced.
Training sequence BotChat WebForm
Llama 3.1 8B (out-of-box) 2.34 3.32
BotChat only 4.265 3.33
WebForm only 2.85 4.215
BotChat→WebForm 2.57 3.405
WebForm→BotChat 2.77 3.41
BotChat+WebForm 4.325 4.405
For both channels, the best performance is achieved when training datasets from both channels are provided simultaneously (BotChat+WebForm). When only a single channel dataset is used for training (BotChat only, WebForm only), the performance for that channel remains similar but slightly lower compared to the BotChat+WebForm configuration. This suggests that despite channel-specific differences, the model benefits from cross-channel learning due to shared underlying principles and evaluation criteria for customer interaction summarization. Moreover, models trained exclusively on one channel (BotChat only, WebForm only) still show slight improvements in the summarization performance for the other channel when compared to the out-of-box Llama 3.1 8B performance, further confirming cross-channel knowledge transfer.
When training data is provided sequentially (BotChat→WebForm, WebForm→BotChat), mean auto-ratings are higher compared to out-of-box models for both channels, demonstrating that the models do gain some knowledge from sequential training. However, these ratings remain notably lower than those achieved through simultaneous training (BotChat+WebForm) or single-channel training (BotChat only, WebForm only). For the initially trained channel (BotChat in BotChat→WebForm, WebForm in WebForm→BotChat), there may have been performance degradation due to catastrophic forgetting [12, 13]. Conversely, the performance on the channel trained subsequently (WebForm in BotChat→WebForm, BotChat in WebForm→BotChat) also remains significantly lower than other training configurations, possibly due to reduced learning capacity following initial fine-tuning with LoRa.
8 Discussion
In this paper, we successfully fine-tuned a smaller student LLM to surpass the performance of a teacher LLM on customer service interaction summarization tasks. Through our proposed "Analyze-Revise-Finetune" approach, we demonstrated that Llama 3.1 8B could outperform GPT-3.5. This clearly illustrates the effectiveness of our pipeline for specialized applications, highlighting its potential for replacing large proprietary models with smaller open-source alternatives. Adopting smaller models can significantly mitigate cost and privacy concerns associated with larger, proprietary LLMs.
We established a detailed evaluation framework tailored specifically for production systems that summarize customer service interactions. This framework ensures a consistent and systematic evaluation of summary quality, allowing for precise identification of the most common errors requiring correction. Additionally, by separating the editor LLM from the teacher LLM and utilizing Llama 3.1 70B for training data revisions, our approach further optimizes both inference cost and data privacy, offering substantial advantages compared to approaches that directly employ large teacher models for revisions.
Although performance gains were observed across different student models using the revised training datasets, the quality of the revisions emerged as an important factor influencing final model performance. Thus, future work should focus on developing more robust and reliable revision processes to further improve the overall effectiveness and dependability of the training data revision step.
9 Conclusion
In this paper, we introduced and validated a fine-tuning pipeline incorporating targeted error-correction designed to enhance the quality of training data generated by large proprietary teacher models for smaller, open-source student models. By integrating detailed error analysis and targeted text revisions, our "Analyze-Revise-Finetune" approach achieves strong summarization performance while significantly reducing costs and addressing privacy concerns associated with proprietary systems.
The results suggest that the Analyze–Revise–Finetune pipeline can be successfully extended to a wide range of LLM-based applications, offering consistent improvements in cost efficiency, output quality, and data privacy.
Acknowledgments
We’d like to thank Jillin Gerten, Vishal Khadake, Mehran Nadjarbashi Noghani, Swapnil Thorat, Ryan Wen, Zonghan Xu, Pooja Daryani, and Megha Korade for their support in this work.
Appendix A Additional Table
Table 5: Summary error types as defined in the summary evaluation scheme
Primary Label Sub Label Description
Content content_missing Error in summary content - 1) missing critical content, 2) not following a logical chronological order, 3) inaccurate representation of original content, or 4) inclusion of completely fabricated content
content_order
content_inaccurate
content_hallucination
content_other Any other content issues
Entities entity_missing_item_number An entity (i.e., item number, item name, order number, return number, case number, username, price, or transaction ID) that is represented in the original content is missing in summary
entity_missing_item_name
entity_missing_order_number
entity_missing_return_number
entity_missing_case_number
entity_missing_username
entity_missing_price
entity_missing_transaction_id
entity_inaccurate_item_number An entity (i.e., item number, item name, order number, return number, case number, username, price, or transaction ID) that is represented in the original content is inaccurately presented in summary
entity_inaccurate_item_name
entity_inaccurate_order_number
entity_inaccurate_return_number
entity_inaccurate_case_number
entity_inaccurate_username
entity_inaccurate_price
entity_inaccurate_transaction_id
entity_inaccurate_other When inaccurate entity type is unclear due to redaction
entity_other Any other entity-related issues
Data Elements data_element_missing A data element (e.g., dates) is missing or inaccurate in summary
data_element_inaccurate
data_element_other Any other issues related to data elements
Customer Type customer_type_missing Customer type is missing or inaccurate in summary. Applicable only to WebForm.
customer_type_inaccurate
customer_type_other Other Issues related to customer type
Unnecessary information unn_content_redundant Summary includes unnecesary information, i.e., 1) redundant or repetitive information, 2) courtesy or plesantry, 3) bot’s response, 4) customer ranting, 5) lack of specific details in the original content, 6) requets for human agent, 7) transfer to next agent, 8) request for email copy of original webform
unn_content_courtesy
unn_content_ebay_response_included
unn_content_customer_rant
unn_content_no_details
unn_content_requests_agent
unn_content_transfer
unn_content_webform_email_copy
unn_other Any other type of unnecessary information
Inferred Sentiment sentiment_inferred_confused Inferred customer sentiment, i.e., confusion or frustration, not in the original content but in summary
sentiment_inferred_frustrated
sentiment_inferred_not_confused Inferred info. of customer not showing sentiment, i.e., confusion, frustration, or complaint, not in the original content but in summary
sentiment_inferred_not_frustrated
sentiment_inferred_no_complaint
sentiment_other Any other sentiment-related issues
Language language_non_english_not_identified Missing the fact that oroginal content is in a non-English
language_translation_inaccurate Inaccuracy in summary due to translation
language_other Any other issues related to language